A body of mass m moving in a circular irbit of radius R about a planet of mass M. At some instant it splits into two equal masses the first mass moves in a cicular orbit of radius R/2 and the other mass in a circular orbit of radius 3R/2 the difference between the final and initial total energies is

Grade:12th pass

1 Answers

FIZA SHOUKATHALI
5 years ago
Ei=-GMm/2R 
Ef=-GM (m/2)/2 (R/2)+ -GM(m/2)/2(3R/2)
    = -2GMm/3R
Ef-Ei= -2GMm/3R - -GMm/2R
        = (-GMm/R )(2/3 - 1/2)       
       = -GMm/6R